OpenStudy (anonymous):

4x – y + z = 6 3x + 2y – z = 0 4x – y – 3z = 2

OpenStudy (anonymous):

4x-y+z=6 -/4x-y-3z=2 4z=4 ----> z=1 2/4x-y=5 3x+2y=1 11x=11 -----> x =1 so y=-2

OpenStudy (anonymous):

But my choices are.. 7x + y = 6 x – y = 6 x + y = 6 7x – y = 6

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

y=3

OpenStudy (texaschic101):

you got an elimination problem.....or substitution....can both of you handle it ?

OpenStudy (anonymous):

oh my god I am making mistakes all the time

OpenStudy (anonymous):

y=3 is not one of my choices..sorry

OpenStudy (anonymous):

y=-1 okey

OpenStudy (texaschic101):

you got 3 equations.....not 2...remember that

OpenStudy (anonymous):

so it is the first one 7x + y = 6 correct!!!!

OpenStudy (texaschic101):

Do you need me to do the problem ?

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

That dude dont no what the heck he is doing and id appreciate an example

OpenStudy (texaschic101):

Hold on...this will take a little while...I will soon post the answer :)

OpenStudy (anonymous):

I have already done that

OpenStudy (anonymous):

the answer is 7x+y=6 dude

OpenStudy (anonymous):

I dont need u to give me the answer "DUDE" i want the explanation and example

OpenStudy (anonymous):

@raebaby420 really? do not get it

OpenStudy (anonymous):

And im about to teache you something too... You dont call girls DUDE its a sign of disrespect. and no i do not get the 4 different answers u gave me, thanks for trying. Bye now

OpenStudy (anonymous):

@raebaby420 i did not say "dude" to you :) to texac..I was trying to help but you are so stupid to understand...I am sorry bye

OpenStudy (anonymous):

but you have great boobies I could not argue that

OpenStudy (anonymous):

Im not stupid, im on here to get help and examples and u sir, are not one to give them OBVIOUSLY. Now please stop causing trouble so i can learn hwo to do this from people who are willing to help and get off here. now leave me alone

OpenStudy (anonymous):

an advice..when you want help do not say the people who are trying to help That dude dont no what the heck he is doing

OpenStudy (anonymous):

you perky bastard

OpenStudy (anonymous):

Excuse me. please get off my thing. and leave so people will help me. This is getting ridiculous

Directrix (directrix):

@SkykhanFalcon From the OpenStudy Code of Conduct Don't be offensive, inappropriate, or creepy! So avoid: •Insulting, berating, or otherwise discouraging others. •Cursing and bypassing the language filters. •Sexual innuendo, or "hitting on people". •Soliciting personal, private, or offline contact information. •Any form of racism, sexism, or any other prejudicial "isms".

OpenStudy (anonymous):

Thank you. the moderators wont do anything but im glad u would say something.

OpenStudy (anonymous):

Im definitely complaining about this site later on. This is ridiculous they have moderators that wont say anything but people who use the site do.

OpenStudy (anonymous):

Oh good!! im still a little confused. i understand how you got that far. but how does it relate to my answer choices?

OpenStudy (anonymous):

Appreciate that..x=1, y=-1, z=1 so the answer 7x+y=6 :)

OpenStudy (saifoo.khan):

@SkykhanFalcon : Like @Directrix stated, we have certain rules on this site. Please read the code of conduct of this site here: http://openstudy.com/code-of-conduct Moreover, be nice to users and no NEVER use inappropriate language nor try to engage in any sort of activity you just took part in. We strictly ask our users to follow the code of conduct. If you have any issues, feel free to write us at: abuse@openstudy.com our team will reply you on asap basis. Thanks.

OpenStudy (anonymous):

If you want to learn something, you can check my messages on top :)

OpenStudy (anonymous):

@raebaby420, I'm sorry, I messed up. But you can use the answer @SkykhanFalcon just put.

OpenStudy (anonymous):

..This is going no where

OpenStudy (anonymous):

@Zelda thank you zelda.. I did the question but she is not gonna believe me :)

OpenStudy (anonymous):

I NEED EXAMPLES

OpenStudy (anonymous):

4x-y+z=6 -/4x-y-3z=2 4z=4 ----> z=1 2/4x-y=5 3x+2y=1 11x=11 -----> x =1 -----> y=-1 then your choices are 7x + y = 6 x – y = 6 x + y = 6 7x – y = 6 If you put the values I found in these equations, you can see the answer as 7x+y=6

OpenStudy (anonymous):

@saifoo.khan good site..you know any site for engineering electromagnetics

OpenStudy (saifoo.khan):

Try http://stackexchange.com/ http://stackoverflow.com/

OpenStudy (texaschic101):

Hey listen up....those answer choices cannot be the answer to this problem...if you finished this problem you would get the answer to all the integers for instance....x = ? or y = ?......the answer choices that are given are not in that form. However, when you add just the first 2 equations...this is what you get... 4x - y + z = 6 3x + 2y - z = 0 -------------add 7x + y = 6 This is not the answer to the entire problem, but it is a step in the problem....is this what you want ? If so, then this is your answer :)